# Forecasting Melting Points in Svalbard, Norway Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

### v1

@@ 4.2 The Statistical Procedures Used
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are of particular concern under Arctic amplification, often requiring practical climate adaptations.
If melting days are so important, one might argue for reformulating the analysis as a classification problem—above or below $0^{\circ}\mathrm{C}$. While melting is indeed a critical threshold, the degree to which that threshold is exceeded also matters. The melting process is highly nonlinear and difficult to characterize, but evidence suggests that melting increases at an increasing rate as temperatures rise (Polashenski, Perovich and Courville, 2012; Pizner et al., 2024).
Adaptive conformal prediction regions provide provably valid coverage when data are exchangeable. Multiple time-series data, however, are typically not exchangeable. With an appropriate fitting algorithm, this limitation can be addressed. Details and supporting pseudocode are presented later, in the grounded context of the forecasting results. The approach is at least unusual, and perhaps even novel.
@@ 5.1 Univariate Plots
As a sanity check, a time series of the 2 p.m. temperatures for 2023 should reflect the expected Arctic seasonal swings. Figure 1 (left panel) shows precisely that. The light-blue irregular line interpolates the daily temperatures, while the smooth black line represents a loess smoother applied to those values. The red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$. Temperatures during the nocturnal months are, on average, colder than during the diurnal months, and transitions between them are gradual. Melting temperatures are common from early June through late September.
@@
The right panel shows a histogram of the 2 p.m. temperatures for 2023 with a generalized extreme value (GEV) distribution overlay. Temperatures range from slightly below $20^{\circ}\mathrm{C}$ to slightly above $10^{\circ}\mathrm{C}$, revealing substantial variability to study. The GEV overlay adds little interpretive value. The histogram is roughly symmetric, although the left tail is somewhat longer than the right. No clear outliers are apparent.
@@ 5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that each time series of daily 2 p.m. temperatures for 2022, 2023, and 2024 ideally represents realizations from the same underlying joint probability distribution—the underlying physics for corresponding months in those three years should be comparable. However, concerns have been raised about possible departures due to Arctic amplification.
Figure 2 provides a visual comparison of the three temperature time series. The jagged lines show the 2 p.m. daily temperatures plotted against day of year, and the dashed horizontal line marks the melting temperature at $0^{\circ}\mathrm{C}$.
@@
Figure 2 shows substantial overlap among the three years. Seasonal trends are captured in a similar manner, both in temperature levels and in timing. The time series move largely in lockstep from early to late summer, when melting temperatures are nearly universal. There appears to be no compelling reason to reject the claim that each series consists of random variables drawn from the same joint probability distribution. In practice, the 2022 and 2024 temperatures serve as promising holdout-sample candidates for evaluating forecasts of the 2023 temperatures.
@@ 5.3 The Quantile GBM Fit of Temperature
With the quantile parameter fixed at $\tau=0.60$, the quantile gradient-boosting procedure in R ran efficiently. The shrinkage value was set to 0.0001, the interaction depth to 6, and the minimum number of observations in a terminal node to 6. These values were chosen to foster slow convergence so that the tails of the temperature distribution would be fitted more accurately. Using 2022 as a holdout sample indicated that approximately 27,000 iterations were appropriate for these data. [^4] With a substantially larger shrinkage value, many fewer iterations might have been adequate. However, the best shrinkage value could not be known in advance, and selectively tuning parameters can undermine later statistical inference (Kuchibhotla, Kolassa and Kuffner, 2022).
@@
Figure 3: Observed versus fitted 2 p.m. temperatures. Black dots are the data, the blue solid line is a loess smooth, and the red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$.
@@
Figure 3 plots the observed 2 p.m. temperatures against their fitted values. The relationship is roughly linear and positive: observed and fitted values tend to increase together. The clustering around the smoothed fitted line is somewhat tighter when the fitted values exceed the melting point, with residuals rarely larger than about $\pm 2^{\circ}\mathrm{C}$. This results in part from the heavier weighting of those values through the target quantile $\tau=0.60$. There is greater sparsity on either side of the temperature range between approximately $5^{\circ}\mathrm{C}$ and $0^{\circ}\mathrm{C}$.
Koenker and Machado (1999) discuss measures of fit for conditional-quantile models derived from the quantile loss function. However, no analogue of the multiple correlation coefficient exists for quantile regression. A suitable measure of fit is introduced later, when conformal prediction regions are presented.
@@
Two points merit emphasis. First, the apparent linearity in Figure 3 says nothing about whether the lagged predictors themselves relate linearly to temperature. In Figure 1, the $x$-axis units are days; in Figure 3, they are degrees Celsius. The former shows day-to-day variation in observed temperatures, whereas the latter shows how observed temperatures vary with their fitted counterparts. Second, lagging the predictors means they generate fitted values two weeks before the future temperatures materialize. This is not true forecasting, since no new unlabeled cases are involved, but it is an important step in that direction. [^5] Fitted values are sometimes called predicted values, which can cause confusion. In this paper, the term fitted values refers to outcomes within the training data, whereas forecasted values refer to predictions for new, unseen cases.
@@
Figure 4: Observed daily 2 p.m. temperatures and fitted values from quantile gradient boosting plotted against day of year (2023). The solid red line marks the melting point at $0^{\circ}\mathrm{C}$.
@@
Figure 4 provides a complementary perspective that connects more directly to the Arctic-amplification context. It re-expresses the results from Figure 3 as a time series, plotting temperature against day of year. As before, black dots denote the observed 2 p.m. temperatures, the jagged blue line shows the fitted values from quantile gradient boosting, and the horizontal red line marks the melting point.
@@
Overall, the fitted values track the observed temperatures closely and capture temperatures above $0^{\circ}\mathrm{C}$ especially well. The several unusually warm days clustering in mid-July—around $10^{\circ}\mathrm{C}$—are of particular concern to climate scientists (Semenov, 2021), and they too are fitted accurately, albeit with slight underestimation. Two weeks in advance, the fitted algorithm anticipates melting temperatures effectively, including those mid-July extremes. The fit might improve further if a larger value of $\tau$ were used, but higher quantiles (e.g., $Q(0.90)$) would rely on far sparser data, potentially leading to instability. As with Figure 3, however, true forecasting still remains to be performed.
@@ 5.4 Predictor Impacts on the Fitted Values
Like all algorithms, quantile gradient boosting is not a model in the traditional statistical sense (Breiman, 2001; Kearns and Roth, 2019). Nevertheless, useful insights about associations among variables can be obtained from variable-importance plots and partial-dependence plots (Friedman, 2002).
@@
The prominence of the day counter and the 14-day temperature lag is unsurprising. Clear seasonal patterns are captured by the counter, and temperature’s gradual evolution over time makes its own lag a strong predictor. The other variables may still capture small or localized temporal effects that, while contributing less to predictive accuracy, could be of scientific or policy interest.
@@ 5.5 Functional Forms
The quantile gradient-boosting algorithm learns associations between each predictor and the response, including the shapes of those relationships. The partial-dependence plots in Figure 6 show the relationships between the day counter and the 14-day lagged temperature with the temperature response variable, each computed with all other predictors fixed at their means—that is, all other predictors are “held constant.” These are the two variables that dominate the fit. The black dots represent the partial-dependence values, and the solid blue lines show a loess smooth of those values.
Figure 6: Partial-dependence plots for the day counter and 14-day lagged temperature. The black dots are the partial-dependence values, and the solid blue line is a loess smooth. Both panels show the relationship with the response variable while all other predictors are fixed at their means.
Consistent with Figure 5, strong relationships are evident. The counter-based fitted values range from approximately $3^{\circ}\mathrm{C}$ to $3^{\circ}\mathrm{C}$, and the lagged-temperature fitted values from about $2.5^{\circ}\mathrm{C}$ to $5^{\circ}\mathrm{C}$. The first plot is roughly symmetric and concave, peaking during the summer months, consistent with the longitudinal pattern shown in Figure 4. The second plot is S-shaped, with its steepest slope beginning immediately after the melting point at $0^{\circ}\mathrm{C}$ is reached and plateauing near $2.5^{\circ}\mathrm{C}$ and $5^{\circ}\mathrm{C}$. The lagged observed temperatures have their strongest association with fitted temperatures just above the melting point. The practical significance of this threshold will be discussed in greater depth below. [^6] The short strings of dots visible in the plot represent rounding artifacts; the values differ slightly.
The remaining four predictors also exhibit nonlinear relationships with the temperature response variable, but because of their comparatively small contributions to the fit, their partial-dependence plots are omitted here in the interest of space.
@@ 6 Forecasts and Estimated Uncertainty
Conformal prediction regions begin with a pre-specified coverage probability. By convention, this probability is denoted $(1-\alpha)$, where $\alpha$ lies between 0 and 0.50. In practice, coverage probabilities range from slightly above 0.50 to nearly 1.0, with values between 0.75 and 0.99 most common. The value of $\alpha$ is determined by subjective considerations shaped by the data and the application.
Once a coverage probability is specified, the data are used to determine the length of each prediction region. A conformal prediction region contains the true forecasted value with probability at least $(1-\alpha)$, conditional on the training data and the performance of the fitted algorithm (Vovk, Gammerman and Shafer, 2005; Vovk et al., 2017; Angelopoulos, Barber and Bates, 2024). Higher coverage yields longer regions, trading confidence for precision; lower coverage shortens the regions at the cost of reduced confidence.
The essential requirement for conformal prediction is that the observations used to construct the regions are exchangeable—that is, the order in which they are realized does not affect their joint distribution. Dependence among the observations is permissible so long as exchangeability holds. For example, random sampling without replacement from a finite population induces dependence but still yields exchangeable samples. [^7] Resampling methods can address lack of exchangeability due to data dependence (Politis, Romano and Wolf, 1999, Chaps. 3–4), but such methods are beyond the scope of this paper.
The longitudinal data from the Longyearbyen weather station exhibit temporal dependence, precluding exchangeability. However, if residuals from a fitted model are indistinguishable from white noise, they may be treated as exchangeable and serve as valid non-conformal scores. In that case, adaptive conformal prediction regions can be computed, allowing forecast-specific intervals of varying length (Romano, Patterson and Candès, 2019). The length of each interval depends on the local uncertainty associated with the corresponding forecast.
Residuals from the 2023 quantile gradient-boosting fit displayed temporal dependence, so exchangeability was not initially achieved. Examination of the residual autocorrelation function suggested a simple AR(1) specification as a special case of an ARIMA model (Box et al., 2015). A single autoregressive coefficient and an intercept were estimated. When applied to the 2023 residuals, the estimated AR(1) coefficient was 0.46 and the intercept 1.44-1.44. Temporal dependence was real but modest. An autocorrelation plot of the AR(1) residuals showed no remaining dependence, and a Ljung–Box test for serial correlation produced a $p$-value of 0.31. The AR(1) model can therefore be regarded as part of the training algorithm that generates white-noise residuals. These residuals serve as acceptable non-conformal scores, restoring exchangeability.
The AR(1) model took the form
$r_{t}=\phi_{0}+\phi_{1}r_{t-1}+\epsilon_{t},$
where $t$ indexes day, $r_{t}$ denotes the 2023 boosting residuals, $\phi_{0}$ and $\phi_{1}$ are the intercept and autoregressive coefficient, and $\epsilon_{t}$ represents the ARIMA innovations. With $\phi_{1}>0$, the variance of $\epsilon_{t}$ is smaller than that of $r_{t}$. The non-conformal scores based on $\epsilon_{t}$ are therefore in the scale of the ARIMA residuals rather than the original temperature scale. Consequently, the innovations were inflated by the factor $1/\sqrt{1-\phi_{1}^{2}}$ to restore the appropriate scale; these inflated residuals are the final non-conformal scores.
The 2024 data served as a pristine holdout sample, representing unlabeled cases for true forecasting. Forecasts were produced using the trained quantile gradient-boosting algorithm and its predict() function in R, with the same predictors used earlier for fitting the 2 p.m. temperatures.
For the adaptive conformal prediction regions, a region is computed for each unlabeled case. The inflated ARIMA non-conformal scores were regressed on all predictors and the corresponding forecasts using quantile random forests (QRF) rather than quantile gradient boosting, because the latter proved numerically unstable. [^8] Quantile gradient boosting (QGB) directly minimizes a global quantile loss to estimate conditional quantiles (Friedman, 2001), whereas quantile regression forests (QRF) use variance-based tree splits and recover quantiles from the empirical distribution of responses within the terminal nodes (Meinshausen, 2006). When $\tau$ is near the extremes, sparsity can hinder QGB’s loss minimization. Because QRF computes quantiles after all trees are grown, such sparsity is avoided. Moreover, once residuals are white noise, the bootstrap sampling in QRF does not distort temporal dependence, since none remains. For each case, the QRF output provided its $Q(0.10)$ and $Q(0.90)$ values as lower and upper bounds, respectively, corresponding to a nominal coverage of at least 0.80.
Formally, with a coverage probability of 0.80, $\alpha=0.20$. Let $t=1,\dots,T$ denote observed cases and $T+1$ index a new case with covariates $\mathbf{x}_{T+1}$. The inflated ARIMA innovations, treated as exchangeable non-conformal scores, define a distribution $C$. With $y_{T+1}$ unknown, the trained boosting algorithm yields the forecast $\hat{y}_{T+1}$ given $\mathbf{x}_{T+1}$. The corresponding lower and upper quantiles are
$\hat{q}_{\alpha/2,\,(T+1)}\bigl(C\bigr)\quad\text{and}\quad\hat{q}_{1-\alpha/2,\,(T+1)}\bigl(C\bigr),$
producing the adaptive conformal prediction region
$\bigl[\hat{q}_{\alpha/2,\,(T+1)}\bigl(C\bigr)+\hat{y}_{T+1},\ \hat{q}_{1-\alpha/2,\,(T+1)}\bigl(C\bigr)+\hat{y}_{T+1}\bigr].$
Full pseudocode appears in the Appendix.
The lengths of the 2024 conformal prediction regions varied widely, consistent with Figure 3. Their average half-width was about $\pm 2.5^{\circ}\mathrm{C}$ for days when the forecasted temperature exceeded 0, and about $\pm 4.5^{\circ}\mathrm{C}$ when it was colder. The longest intervals reached roughly $\pm 4.5^{\circ}\mathrm{C}$ for warm days and $\pm 8^{\circ}\mathrm{C}$ for cold ones.
Interpreting these uncertainty estimates may be difficult for stakeholders. Perhaps more relevant is the likelihood that the actual future future temperature will exceed $0^{\circ}\mathrm{C}$ given a particular forecast. For forecasted temperatures above 0, these represent true positives; for forecasted temperatures at or below 0, false negatives.

### v2

@@ 4.2 Some Details
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when the loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are a particular concern because practical climate adaptations at a local level are often needed.
If melting days are so important, one might argue for reformulating the analysis as a classification problem—above or below $0^{\circ}\mathrm{C}$. While melting is indeed a critical threshold, the degree to which that threshold is exceeded also matters. The melting process is highly nonlinear and difficult to characterize, but current evidence suggests that melting can increase at an increasing rate as temperatures rise (Polashenski, Perovich and Courville, 2012; Pizner et al., 2024).
4.3 Some Details
[…]
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when the loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are a particular concern because practical climate adaptations at a local level are often needed.
If melting days are so important, one might argue for reformulating the analysis as a classification problem—above or below $0^{\circ}\mathrm{C}$. While melting is indeed a critical threshold, the degree to which that threshold is exceeded also matters. The melting process is highly nonlinear and difficult to characterize, but current evidence suggests that melting can increase at an increasing rate as temperatures rise (Polashenski, Perovich and Courville, 2012; Pizner et al., 2024).
@@ 5.1 Univariate Plots
A time series of the 2 p.m. temperatures for 2023 should reflect the expected Arctic seasonal swings. Figure 1 (left panel) shows precisely that. The light-blue irregular line interpolates the daily temperatures, while the smooth black line represents a loess smoother applied to those values. The red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$. Temperatures during the nocturnal months are, on average, colder than during the diurnal months, and transitions between them are gradual. Melting temperatures are common from early June through late September.
@@
The right panel shows a histogram of the 2 p.m. temperatures for 2023 with a generalized extreme value (GEV) distribution overlay. Temperatures range from slightly below $20^{\circ}\mathrm{C}$ to slightly above $10^{\circ}\mathrm{C}$, revealing substantial variability. The GEV overlay adds little interpretive value. The histogram is roughly symmetric, although the left tail is somewhat longer than the right. No clear outliers are apparent.
@@ 5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that each time series of daily 2 p.m. temperatures for 2022, 2023, and 2024 represents realizations from the same underlying joint probability distribution; the underlying physics for corresponding months in those three years should be comparable. However, there could be aberrations caused by Arctic amplification.
Figure 2 provides a visual comparison of the three temperature time series. The jagged lines show the 2 p.m. daily temperatures plotted against day of year. The dashed horizontal line marks the melting temperature at $0^{\circ}\mathrm{C}$.
@@
Figure 2 shows substantial overlap across the three years. Seasonal trends are captured similarly, both in temperature values and in timing. The time series move largely in lockstep from early to late summer, during which melting temperatures are nearly universal. There appears to be no compelling reason to reject the claim that each series consists of random variables drawn from the same joint probability distribution. In practice, the 2022 and 2024 temperatures serve as promising holdout samples for evaluating forecasts of the 2023 temperatures.
@@ 5.3 The Quantile GBM Fit of Temperature
Some changes in the predictor variables were required. Visibility and atmospheric pressure were dropped because of excessive amounts of missing data. Wind direction (reported in degrees) was transformed into its sine and cosine components after converting degrees to radians. This representation captures the circular nature of wind direction, for which 0° and 360° denote the same physical direction. [^5] Wind direction was recorded in degrees but is intrinsically circular, with $0^{\circ}$ and $360^{\circ}$ representing the same direction. Each direction $\theta$ was therefore converted to radians and encoded as $\sin(\theta)$ and $\cos(\theta)$. Using both components is necessary because neither $\sin(\theta)$ nor $\cos(\theta)$ alone uniquely identifies a direction (e.g., $\sin(30^{\circ})=\sin(150^{\circ})$), whereas the pair $(\cos\theta,\sin\theta)$ provides a unique and smooth representation of all possible directions on the unit circle.
@@
With the quantile parameter fixed at $\tau=0.60$, the quantile gradient-boosting procedure in R ran efficiently. The shrinkage value was set to 0.0001, the interaction depth to 6, and the minimum number of observations in a terminal node to 6. These values were chosen to foster slow convergence so that the right tail of the temperature distribution would be fitted more accurately. Using 2022 as a holdout sample to protect against overfitting indicated that approximately 27,000 iterations were appropriate for these data. [^6] With a substantially larger shrinkage value, many fewer iterations might have been adequate. However, the best shrinkage value could not be known in advance, and empirical tuning encourages “cherry picking” that can undermine later statistical inference (Kuchibhotla, Kolassa and Kuffner, 2022).
@@
Figure 3: Observed versus fitted 2 p.m. temperatures. Black dots are the data, the blue solid line is a loess smooth serving as a visual aid, and the red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$.
@@
Figure 3 plots the observed 2 p.m. temperatures against their fitted values. The relationship is roughly linear and positive; observed and fitted values tend to increase together. Clustering around the smoothed fitted line is somewhat tighter when the fitted values exceed the melting point, with residuals rarely larger than about $\pm 2^{\circ}\mathrm{C}$. This results in part from the larger cost for underfitting imposed by the target quantile $\tau=0.60$. There is greater sparsity on either side of the temperature range between approximately $5^{\circ}\mathrm{C}$ and $0^{\circ}\mathrm{C}$. In short, the fit quality produced by quantile gradient boosting is encouraging. [^7] Koenker and Machado (1999) discuss measures of fit for conditional-quantile models derived from the quantile loss function. However, apparently no analogue of the multiple correlation coefficient exists for quantile regression. A suitable measure of fit is introduced later, when conformal prediction regions are presented.
@@
Two points merit emphasis. First, the apparent linearity in Figure 3 says nothing about whether the lagged predictors themselves relate linearly to temperature. In Figure 1 left panel, the $x$-axis units are days; in Figure 3, they are degrees Celsius. The former shows day-to-day variation in observed temperatures, whereas the latter shows how observed temperatures vary with their fitted counterparts. Second, lagging the predictors means they generate fitted values two weeks before the future temperatures are realized. This is not yet true forecasting, because no new unlabeled cases are involved, but it is an important step in that direction. [^8] Fitted values are sometimes called predicted values, which can cause confusion. In this paper, fitted values refer to outcomes within the training or calibration data, whereas forecasted values refer to predictions for new, unlabeled cases.
@@
Figure 4: Observed and Fitted 2 p.m. temperatures from quantile gradient boosting plotted against day of year (2023). Black dots show the observed temperatures over time, the jagged blue line is an interpolation of the fitted values, and the solid red line marks the melting point at $0^{\circ}\mathrm{C}$.
@@
Figure 4 provides a complementary perspective that connects more directly to the Arctic-amplification context. It reproduces the observed temperatures and fitted values from Figure 3 as time series data, plotted against day of the year. As before, black dots denote the observed 2 p.m. temperatures, the jagged blue line in an interpolation of the fitted values, and the horizontal red line marks the melting point.
@@
Overall, the fitted values track the observed temperatures rather closely and capture temperatures above $0^{\circ}\mathrm{C}$ especially well. The unusually warm days clustering in mid-July—around $10^{\circ}\mathrm{C}$—are of particular concern to climate scientists (Semenov, 2021), and they too are fitted accurately, albeit with slight underestimation. Two weeks in advance, the fitted algorithm anticipates melting temperatures effectively, including those mid-July extremes. The fit might improve further if a larger value of $\tau$ were used, but higher quantiles (e.g., $Q(0.90)$) would rely on far sparser data, potentially leading to instability. As with Figure 3, however, true forecasting remains to be performed.
@@ 5.4 Predictor Impacts on the Fitted Values
Like all algorithms, quantile gradient boosting is not a model in the traditional statistical sense (Breiman, 2001; Kearns and Roth, 2019). Nevertheless, useful insights about associations in the data can be obtained from variable-importance plots and partial-dependence plots (Friedman, 2002).
@@
The prominence of the day counter and the 14-day temperature lag is unsurprising. Seasonal patterns are captured by the counter, and the temperature’s gradual evolution over days makes its own lag a strong predictor. The other variables may still capture small or localized temporal effects that, while contributing less to fitting performance, are nonetheless systematically related to the melting temperatures.
@@ 5.5 Functional Forms
The quantile gradient-boosting algorithm learns associations between each predictor and the response, including the shapes of those relationships. The partial-dependence plots in Figure 6 show the relationships between the day counter in the left and the 14-day lagged temperature on the right with the 2 p.m. temperature response, each computed with all other predictors fixed at their means. These are the two variables that dominate the fit. The black dots represent the partial-dependence values, and the solid blue lines show a loess smooth as a visual aid.
Figure 6: Partial-dependence plots for the day counter and the 14-day lagged 2 p.m. temperature. The black dots are the partial-dependence values, and the solid blue line is a loess smooth serving as a visual aid. Other predictors are fixed at their means.
Consistent with Figure 5, strong relationships are evident. A look again at Figure 4 provides context. The counter-based fitted values range from approximately $3^{\circ}\mathrm{C}$ to $3^{\circ}\mathrm{C}$, and the lagged-temperature fitted values from about $2.5^{\circ}\mathrm{C}$ to $5^{\circ}\mathrm{C}$. The first plot is roughly symmetric and concave, peaking during the summer months, consistent with the longitudinal pattern shown in Figure 4. The second plot is S-shaped, with its steepest slope beginning immediately after the melting point at $0^{\circ}\mathrm{C}$ and plateauing near $2.5^{\circ}\mathrm{C}$ at the left and $5^{\circ}\mathrm{C}$ at the right. The lagged observed temperatures have their strongest association with fitted temperatures just above the melting point. The practical significance of this threshold will be discussed in greater depth below. [^9] The short horizontal strings of dots visible in the plot represent rounding artifacts; the values differ slightly. The remaining four predictors also exhibit nonlinear relationships with the temperature response, but because of their comparatively small contributions to the fit, their partial-dependence plots are omitted in the interest of space.
@@ 6 Forecasts and Estimated Uncertainty
Addressing uncertainty in a principled way has been anticipated in the research design and algorithmic procedures employed. The 2024 data serve as a pristine holdout sample of test data that can represent unlabeled cases used in genuine forecasting. [^10] The 2024 data have labels that are useful for didactic purposes and for evaluation of forecasting accuracy. But those labels have no role in the forecasts themselves and would not be known in an operational setting. Forecasts for the 2024 data can be produced using the previously fitted quantile gradient-boosting algorithm and its predict() function in R. From the perspective of the trained algorithm, the 2024 lagged predictors constitute new, unlabeled data.
@@
Obtaining the adaptive conformal prediction regions is more involved. All conformal prediction regions begin with a pre-specified coverage probability. By convention, this probability is denoted by $(1-\alpha)$, where $\alpha$ lies between 0 and 0.50. In practice, coverage probabilities range from slightly above 0.50 to nearly 1.0, with values between 0.75 and 0.95 most common. The value of $\alpha$ is determined by subjective considerations shaped by the data and the application. For the conformal approach used with the Svalbard data, the coverage is $1-\alpha=0.80$.
Larger coverage probabilities lead to larger conformal prediction regions. If the goal is to be highly confident that a forecast will be found within a particular prediction region, it makes sense that the length of the prediction region should be long. If the goal is to reduce the length of the prediction region in service of greater precision, confidence that the forecast will be found within that region must decline. Conformal prediction regions formalize this trade-off between coverage and precision. Stakeholders might legitimately prefer a different value for coverage, and within limits imposed by the data, another coverage value is easily implemented. [^11] Ideally, a future decision about a coverage probability would be made with new data. If a new coverage probability is selected for the 2024 data after seeing the results, new complications are introduced and the conformal methods would need some modest changes (Sarkar and Kuchibhotla, 2023).
Conformal prediction regions are constructed from “nonconformal scores.” These scores can be derived directly from the calibration-data residuals when the calibration data are used as the input into the previously trained boosting algorithm (i.e., predict() in R). Because temperatures from the Svalbard calibration data have strong temporal dependence and the boosting residuals do as well, the residuals were whitened using a first-order autoregressive model (AR(1)). This removes the short-range serial dependence while preserving the level and variability required for valid nonconformal scores. [^12] These steps were guided by estimated autocorrelation functions and Box–Ljung tests determining whether the AR(1) model produced white noise residuals.
Exchangeability is required for valid prediction regions (Vovk, Gammerman and Shafer, 2005; Vovk et al., 2017; Angelopoulos, Barber and Bates, 2024). One implication of exchangeability is that it does not matter, for the theoretical guarantees, that a test observation is realized after all of the calibration observations. A formal summary of conformal prediction is provided in Appendix A along with some computation details. Angelopoulos and Bates (2023) provide an accessible introduction to conformal prediction.
Figure 7 illustrates how conformal prediction regions work. The figure highlights important features of the Svalbard data but is not meant to be a reproduction of those data. On the horizontal axis is the forecasted temperature. On the vertical axis is the observed future temperature that would, in practice, be unknown when a forecast is made. The solid blue line conveys how the forecasts and future temperatures are systematically related. The dashed red lines are the upper and lower bounds of the adaptive prediction region for each forecast. If $1-\alpha=0.80$, the prediction region for any forecast contains the correct future temperature with a probability of at least 0.80.
Figure 7: Simulation of adaptive conformal prediction regions from quantile gradient boosting with $\tau=0.80$. The horizontal axis is the forecasted temperature. The vertical axis is the observed future temperature. The solid blue line shows how forecasts and true future temperatures are systematically related. The dashed red lines show the upper and lower bounds of the adaptive prediction region, and the two vertical dotted magenta lines mark the prediction regions when the forecasted temperature happens to be $10^{\circ}\mathrm{C}$ or $5^{\circ}\mathrm{C}$.
To help fix these ideas, consider two illustrative forecasts, one at $10^{\circ}\mathrm{C}$ and the other at $5^{\circ}\mathrm{C}$. Several features of their prediction regions are apparent and apply to all forecasted temperatures in the figure. Forecasts are not necessarily in the middle of their prediction regions, which means that the regions need not be symmetric around the forecast. In addition, because for the quantile gradient-boosting algorithm $\tau$ was set to 0.60, the forecasts tend to fall above each region’s median. Finally, the lengths of the adaptive prediction regions grow shorter as the fit improves (Romano, Patterson and Candès, 2019). In Figure 7, forecasts above $0^{\circ}\mathrm{C}$ fit the observed temperatures better and have shorter region lengths. In the Svalbard setting, forecasts would be produced one day at a time, but Figure 7 is meant to convey the results after many forecasts are made. The figure also makes plain that there is no requirement that the forecasts be correct.
Although the nominal tail probabilities defining the lower and upper limits are symmetric (e.g., $\alpha/2$ and $1-\alpha/2$), the resulting prediction regions need not be. The asymmetry arises because the values of the upper and lower bounds are determined by the empirical distribution of the nonconformal scores themselves. Their ranks determine the quantiles, but the nonconformal scores need not be equally spaced from one another. For example, if the large positive scores are less dispersed than the large negative scores, the prediction region above the forecasts will be shorter than the prediction region below the forecast.
6.1 Uncertainty Results
The lengths of the 2024 conformal prediction regions vary substantially, consistent with Figure 3. For days in which the forecasted 2 p.m. temperature exceeds 0, the average half-width is about $\pm 2.6^{\circ}\mathrm{C}$, and the half-widths range from about $1.4^{\circ}\mathrm{C}$ to $6.2^{\circ}\mathrm{C}$. For days in which the forecasted 2 p.m. temperature does not exceed 0, the average half-width is about $\pm 4.7^{\circ}\mathrm{C}$, and the half-widths range from about $2.0^{\circ}\mathrm{C}$ to $8.3^{\circ}\mathrm{C}$. Performance is better for the warmer diurnal temperatures because $\tau>0.50$.
Stakeholders might not find much comfort in these summary statistics. Even if the mean half-widths promise some guidance for policy and practice, the ranges of the half-widths counsel caution. Perhaps more useful information could be found in estimates of the distributions of the future observed temperatures for each forecasted temperature. In principle, there would be more information, but the sample sizes would be too small, even over several years of data, to support such estimates properly.
In the spirit of conditional distributions of future temperatures, a more modest approach might work. The forecasts can be binned to increase the sample sizes, and rather than estimating full conditional distributions, estimate how likely a future temperature of $0^{\circ}\mathrm{C}$ is exceeded. Treating melting as a “positive”, when a forecasted temperature is above $0^{\circ}\mathrm{C}$ and the future temperature $0^{\circ}\mathrm{C}$, a true positive results. When the forecasted temperature is $\leq 0^{\circ}\mathrm{C}$ and the future temperature is $0^{\circ}\mathrm{C}$, one has a false negative. Ideally, true positives are far more common than false negatives.